A tennis ball of mass 56 gm is projected vertically upward with average velocity 48.1 m/s. Calculate the kinetic energy after 5 seconds
1
Expert's answer
2020-07-09T10:33:33-0400

"KE=\frac{mv^2}{2}"


"v_f=v_0-gt"


When "v_f=0" "t=4.903s"


So,


"5-4.903=0.097s"


In this case


"v_f=v_0+gt=0+gt=gt=9.81\cdot0.097=0.95m/s"


Finally


"KE=\frac{mv^2}{2}=\frac{0.056\cdot 0.95^2}{2}\approx0.03J"